OpenStudy (anonymous):

Simultaneous equation 0=2a^2+2b 8=4a^2+4b X= Y=

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

can u solve it as i tried and codnt

OpenStudy (anonymous):

So you're trying to find a and b that satisfy the above equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh lol i wrote x and y instead of a and b

OpenStudy (anonymous):

It doesn't seem to work. No values of a and b works

OpenStudy (anonymous):

crap i must have somin wrong let me tell the word that i changet into the equation

OpenStudy (anonymous):

the curve y=ax^2+bx passes through (2,0) and (4,8) find a and b

OpenStudy (anonymous):

0 = 4a +2b 8 =16a + 4b 0 = 4a +2b 0 = 8a + 4b 8=16a +4b - 0 = 8a+4b 8 = 8a a =1 b = -2